Suppose you bought 900 shares of stock at an initial price of $44 per share. The stock paid a dividend of $0.42 per share during the following year, and the share price at the end of the year was $39.

a. Compute your total dollar return on this investment. (A negative value should be indicated by a minus sign.)

b. What is the capital gains yield? (A negative value should be indicated by a minus sign. Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

c. What is the dividend yield? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

d. What is the total rate of return on the investment? (A negative value should be indicated by a minus sign. Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

Next Visit question map

Question1of10Total1 of
Answer #1

Given ,

P0 = $ 44             P1 = $ 39        D1 = $ 0.42

a) Total dollar return on investments is the total of dividend earned and the amount of appreciation/ depreciation in the stock price.

D1 + ( P1 - P0 )

= 0.42 + ( 39 - 44 )

= - 4.58 $

Total return = -$ 4.58 * 900 shares = - $ 4,122

b) Capital gain is the difference between price at the end of the year and the price at the beginning of the year. Capital gain yield is -

= ( P1 - P0 ) / P0

= ( 39 - 44 ) / 44

= - 0.1136 = -11.36 %

c) Dividend yield = D1 / P0

= 0.42 / 44

= 0.0095 i.e., 0.95%

d) Total rate of return on investment is

= [D1 + ( P1 - P0 )] / P0

= [ 0.42 + (39 - 44) ] / 44

= - 4.58 / 44

= - 0.1040    i.e., - 10.4 %